# An Information Theoretic Perspective on Conformal Prediction

### v2

@@ 1 Introduction
In this work, we take a closer look at conformal prediction through the lens of information theory (IT), establishing a connection between conformal prediction and the underlying intrinsic uncertainty of the data-generating process, as captured by the conditional entropy $H(Y|X)$ of the target variable $Y$ given the inputs $X$. We prove conformal prediction can be used to bound $H(Y|X)$ from above in three different ways: one derived from the data processing inequality, which we dub DPI bound, and two coming from a variation of Fano’s inequality [18], a model agnostic one, which we call (simple) Fano bound, and another informed by the predictive model itself, to which we refer as model-based Fano bound. To the best of our knowledge, these bounds represent the first bridge connecting information theory and conformal prediction, which we hope will bring new tools to both fields. We already present two such tools in this paper: (i) we show our upper bounds serve as principled training objectives to learn machine learning models that are more amenable to SCP, and (ii) we advance a systematic way to incorporate side information into the construction of prediction sets. We empirically validate that both these applications of our theoretical results lead to better predictive efficiency, i.e. narrower and, consequently, more informative prediction sets.
@@ 2 Background
In this section, we present the needed background on conformal prediction and list decoding [17, 66], an area of information theory that, as we show, is closely related to CP and especially useful in deriving our main results. We start with the necessary notation. As usual, we denote random variables in uppercase letters and their realization in lowercase, e.g., $X=x$. We reserve calligraphic letters, e.g. ${\mathcal{X}}$, for sets and use $P_{X},Q_{X},\dots$ to denote probability measures on the space ${\mathcal{X}}$. To simplify the notation, we use $P,Q,\dots$ when the underlying space is clear. For example, given a probability measure $Q_{XY}$, the probability of the event $\{(X,Y):Y\in{\mathcal{C}}(X)\}$ is denoted as $Q(Y\in{\mathcal{C}}(X))$.
@@ 2.1 Conformal Prediction
The first step in SCP is to define a nonconformity score function $s_{f}:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow{\mathbb{R}}$, which is itself a function of model $f$ and captures the magnitude of the prediction error at a given data point; the higher the score $s_{f}(x,y)$, the higher the disagreement between input $x$ and prediction $y$. At calibration time, evaluate the score function at every $(X_{i},Y_{i})\in{\mathcal{D}}_{cal}$ to get a collection of scores $\{S_{i}=s_{f}(X_{i},Y_{i})\}_{i=1}^{n}$, and at test time, we construct prediction set ${\mathcal{C}}(X_{test})$ as
@@
where $\text{Quantile}(1-\alpha;\{S_{i}\}_{i=1}^{n})\}$ is the level $1-\alpha$ quantile of the empirical distribution defined by $\{S_{i}\}_{i=1}^{n}$. The central result in conformal prediction, which we restate below for completeness, proves that prediction sets thus constructed achieve marginal valid coverage, i.e., satisfy (1).
@@ 2.2 Conformal Prediction as List Decoding
In a nutshell, we can see conformal prediction as defining a map from an input $x$ to a set of candidates in the target set ${\mathcal{Y}}$. It turns out that the conformal prediction framework is equivalent to (variable-size) list decoding, an error-recovery model going back to the works of Elias [17] and Wozencraft [66] in communication theory—we review some of these results in Appendix C.2. In particular, consider the mapping from the true label $y$ to the input $x$ as a noisy communication channel $p(x|y)$. The goal of an error-correcting code is then to decode the input $x$ and recover the one true label $y$. List decoding generalizes this idea, allowing the decoder to return a set of outcomes (a list) instead of a point-wise prediction. If the correct solution is not part of the set output by the decoder, an error is declared. Although conformal prediction and list decoding were developed for different purposes, namely uncertainty quantification and error correction, it is easy to see that, if we allow for variable-size lists, the list decoding problem for the channel $p(x|y)$ as described above is equivalent to the conformal prediction problem.
@@ 3 Information Theory Applied to Conformal Prediction
In this section, we develop our main results, which link information theory and conformal prediction. Concretely, we provide three novel upper bounds on the conditional entropy $H(Y|X)$: one coming from the data processing inequality and two derived in a similar way to Fano’s inequality. We defer the proofs to Appendix D for the sake conciseness, but in broad strokes, our results come from relating the bounds on the error probability provided by these information-theoretic inequalities (typically in the context of error-correcting codes) to the guarantees provided by CP in Theorem 2.1.
@@ 3.1 Data Processing Inequality for Conformal Prediction
We start by using the classical data processing inequality (DPI) in the context of conformal prediction. Specifically, we focus on the DPI for $f$-divergences, which discuss thoroughly in Appendix C.1. In brief, for a convex function $f$ with $f(1)=0$, the $f$-divergence between two probability measures $P$ and $Q$ is defined as (see [56])
@@
In particular, with $f(x)=x\log x$ we recover the familiar notion of KL-divergence [25]. The DPI for $f$-divergences states that for any two probability measures $P_{X}$ and $Q_{X}$ defined on a space ${\mathcal{X}}$, and any stochastic map $W_{Y|X}$, which maps $(P_{X},Q_{X})$ to $(P_{Y},Q_{Y})$, we have
@@
Note that the entropy term $H(Y|X)$ is computed using the measure $P$. We relegate the proof to Appendix D.1. In the bound in (3), the term $Q(Y\in{\mathcal{C}}(X))$ appears inside a log, so an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(X))$ would result in a lower bound and would be biased. We can provide an upper confidence bound on this estimate using the empirical Bernstein inequality [39], and use that instead. Based on the empirical Bernstein inequality, with probability $1-\delta$, we have
@@ 4 Conformal Training
Although split conformal prediction is applicable to any pretrained ML model as a post-processing step, the overall performance of any CP method (commonly measured by its inefficiency) is highly dependent on the underlying model itself. Therefore, previous works have proposed to take CP into account already during model training and directly optimize for low predictive inefficiency [8, 59]. We use the term conformal training to refer to such approaches in general (see Appendix F for an overview of the topic). In particular, we focus on ConfTr [59] since it generalizes and outperforms [8]. In ConfTr [59] each training batch ${\mathcal{B}}$ is split into calibration ${\mathcal{B}}_{cal}$ and test ${\mathcal{B}}_{test}$ halves to simulate the SCP process (see Section 2) for each gradient update of model $f$ and minimize the following size loss
@@
where ${\mathcal{C}}_{f}(x)$ is constructed using the statistics of ${\mathcal{B}}_{cal}$. Here we use the notation ${\mathcal{C}}_{f}(x)$ to make explicit the dependence of the prediction set on the model $f$. Still, SCP involves step functions and Stutz et al. [59] introduce a couple of relaxations to recover a differentiable objective: (i) the computation of quantiles is relaxed via differentiable sorting operators [16, 9, 48]; (ii) the thresholding operation in the construction of prediction sets in (2) is replaced by smooth assignments of labels to prediction sets via the logistic sigmoid.
Our upper bounds on $H(Y|X)$, namely DPI, MB Fano and simple Fano, presented in the previous section can be made differentiable in the same way and thus can also serve as proper loss functions for conformal training. The motivation for doing so is twofold. First, the conditional entropy $H(Y|X)$ captures the underlying uncertainty of the task, or equivalently, the uncertainty under the true labelling distribution $P_{Y|X}$. Thus, by minimizing these upper bounds, we can hope to push the model $f$ closer to the true distribution, which is known to achieve minimal inefficiency [63]. Interestingly, the cross-entropy loss also bounds $H(Y|X)$ from above and thus can be motivated as a conformal training objective from the same angle. In that regard, the DPI bound from Proposition 3.1 is particularly advantageous as it is provably tighter than the cross-entropy (see Appendix D.1).
Second, we can connect the simple Fano bound from Corollary 3.1 to the size loss (7) from [59]. In Appendix F.1, we show that via Jensen’s inequality and $\log(|Y|{-}|C(X)|)\leq\log|Y|$ the bound in Corollary 3.1 can be further upper bounded as
@@ 5 Side Information
where $Q_{Z|X,Y}$ is an auxiliary model that predicts the side information, given the input and target variables. Such a model could be learned separately from the main model $Q_{Y|X}$ given access to a dataset ${\mathcal{D}}_{side}=\{(x_{i},y_{i},z_{i})\}$. Therefore, we can now calibrate with CP by taking into account the side information and then, at test time, given access to the input and side information, we can use the appropriate probabilities $Q_{Y|X,Z}$ to construct the prediction sets. Intuitively, the prediction sets from such a procedure should be smaller than the prediction sets obtained from using $Q_{Y|X}$ directly. It should be noted that, in the case of side information not being available, we can marginalize $Q_{Y,Z|X}$ over $Z$, which, by construction, falls back to the original model $Q_{Y|X}$. If the availability pattern of $Z$ is consistent between the calibration and test sets, the conformal prediction guarantees still hold by defining the model as
@@
With this addition to the split conformal prediction tool set, if new (side) information is made available at test time, we can properly incorporate it into the CP pipeline without having to retrain the main classifier. One needs only train an auxiliary classifier $Q_{Z|X,Y}$, which might be much simpler than $Q_{Y|X}$ (in our experiments $Q_{Z|X,Y}$ is given by a single linear layer) and require much less data. One notable example of side information arises naturally in the distributed setting, which we discuss next.
@@ 5.1 The Distributed Learning Setting
This gives us an upper bound on the entropy of the entire population that decomposes into a sum of local functions, one for each client, only requiring local information. As a result, we can easily carry out conformal training for $Q_{Y|X}$ by optimizing this upper bound in the federated setting with, e.g., federated averaging [40]. At test time, we can take the device ID $z$ as side information. To this end, we can either train in parallel a model $Q_{Z|X,Y}$ and use (9) with the global model $Q_{Y|X}$ at test time in order to compute an adjusted model $Q_{Y|X,Z}$, or we can obtain $Q_{Y|X,Z}$ by fine-tuning the global model $Q_{Y|X}$ with local data.
@@ 6 Related Work
Conformal prediction, a powerful framework for uncertainty quantification developed by Vovk and collaborators [62, 58] has recently witnessed a wide adoption in many fields, e.g., healthcare [46, 3, 37, 36] and finance [65, 6]. The marriage of conformal prediction and machine learning has been especially fruitful. Since the seminal work by Vovk et al. [62], many extensions and applications have been proposed, covering topics such as survival analysis [10], treatment effect evaluation [30], classification [20, 4] and regression [52] settings, risk control [5, 7], and covariate shift [60].
To our knowledge, our work represents the first attempt to bridge conformal prediction and information theory. Among other things, this allows us to build on the conformal training ideas of Bellotti [8] and Stutz et al. [59], deriving principled learning objectives that generalize their approaches, dispense with some their hyperparameters and result in more efficient prediction sets. Further, we empirically show that our conformal training objectives provide a strong enough learning signal to train complex architectures from scratch, with strong results on ResNet-34 and ResNet-50 [21] fitted on CIFAR10 and CIFAR100, respectively. In contrast, the previous state-of-the-art approach, ConfTr, struggles in those settings (see experiments in Section 7) and required pretrained models for consistent results [59]. Further, our information-theoretic interpretation of CP provides a new simple and effective mechanism to leverage side information in split conformal prediction. We are unaware of any other approaches to treat side information inside the conformal prediction framework in the literature.
On the information theory side, the notion of $f$-divergence and related inequalities have appeared in many different works. The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s, as in, for instance, [2, 14, 42]. A key $f$-divergence inequality is the data processing inequality—see [56, 55] for an extensive survey. It provides a unified way of obtaining many classical and new results, including Fano’s inequality [18]. The tightness of the data processing inequalities is discussed in terms of Bregman’s divergence in [12, 32] and in terms of $\chi^{2}$-divergence in [55]. List decoding, which is closely connected to CP, was introduced in context communication design by Elias [17] and Wozencraft [66]. A generalization of Fano’s inequality to list decoding was given in [15] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [51]. Variable-size list decoding was discussed in [55], using ideas first introduced in [50] and [33]. A selection of relevant inequalities for list decoding can be found in [55].

### v3

@@ 1 Introduction
In this work, we take a closer look at conformal prediction through the lens of information theory (IT), establishing a connection between conformal prediction and the underlying intrinsic uncertainty of the data-generating process, as captured by the conditional entropy $H(Y|X)=-E_{P_{XY}}[\log P_{Y|X}]$ of the target variable $Y$ given the inputs $X$. We prove conformal prediction can be used to bound $H(Y|X)$ from above in three different ways: one derived from the data processing inequality, which we dub DPI bound, and two coming from a variation of Fano’s inequality [19], a model agnostic one, which we call (simple) Fano bound, and another informed by the predictive model itself, to which we refer as model-based Fano bound. To the best of our knowledge, these bounds represent the first bridge connecting information theory and conformal prediction, which we hope will bring new tools to both fields. We already present two such tools in this paper: (i) we show our upper bounds serve as principled training objectives to learn classifiers that are more amenable to SCP, and (ii) we advance a systematic way to incorporate side information into the construction of prediction sets. In a number of classification tasks, we empirically validate that both these applications of our theoretical results lead to better predictive efficiency, i.e., narrower and, consequently, more informative prediction sets.
@@ 2 Background
In this section, we present the needed background on conformal prediction and list decoding [18, 68, 22], an area of information theory that, as we show, is closely related to CP and especially useful in deriving our main results. We start with the necessary notation. As usual, we denote random variables in uppercase letters and their realization in lowercase, e.g., $X=x$. We reserve calligraphic letters, e.g. ${\mathcal{X}}$, for sets and use $P_{X},Q_{X},\dots$ to denote probability measures on the space ${\mathcal{X}}$. To simplify the notation, we use $P,Q,\dots$ when the underlying space is clear. For example, given a probability measure $Q_{XY}$, the probability of the event $\{(X,Y):Y\in{\mathcal{C}}(X)\}$ is denoted as $Q(Y\in{\mathcal{C}}(X))$.
@@ 2.1 Conformal Prediction
The first step in SCP is to define a nonconformity score function $s_{f}:{\mathcal{X}}\times{\mathcal{Y}}\rightarrow{\mathbb{R}}$, which is itself a function of model $f$ and captures the magnitude of the prediction error at a given data point; the higher the score $s_{f}(x,y)$, the higher the disagreement between input $x$ and prediction $y$. At calibration time, we evaluate the score function at every $(X_{i},Y_{i})\in{\mathcal{D}}_{cal}$ to get a collection of scores $\{S_{i}=s_{f}(X_{i},Y_{i})\}_{i=1}^{n}$, and at test time, we construct prediction set ${\mathcal{C}}(X_{test})$ as
@@
where $\text{Quantile}(1-\alpha;\{S_{i}\}_{i=1}^{n})$ is the level $1-\alpha$ quantile of the empirical distribution defined by $\{S_{i}\}_{i=1}^{n}$. The central result in conformal prediction, which we restate below for completeness, proves that prediction sets thus constructed achieve marginal valid coverage, i.e., satisfy (1).
@@ 2.2 Conformal Prediction as List Decoding
In a nutshell, we can see conformal prediction as defining a map from an input $x$ to a set of candidates in the target set ${\mathcal{Y}}$. It turns out that the conformal prediction framework is equivalent to (variable-size) list decoding, an error-recovery model going back to the works of Elias [18] and Wozencraft [68] in communication theory—we review some of these results in Appendix C.2. In particular, consider the mapping from the true label $y$ to the input $x$ as a noisy communication channel $p(x|y)$. The goal of an error-correcting code is then to decode the input $x$ and recover the one true label $y$. List decoding generalizes this idea, allowing the decoder to return a set of outcomes (a list) instead of a pointwise prediction. If the correct solution is not part of the set output by the decoder, an error is declared. Although conformal prediction and list decoding were developed for different purposes, namely uncertainty quantification and error correction, it is easy to see that, if we allow for variable-size lists, the list decoding problem for the channel $p(x|y)$ as described above is equivalent to the conformal prediction problem.
@@ 3 Information Theory Applied to Conformal Prediction
In this section, we develop our main results, which link information theory and conformal prediction. Concretely, we provide three novel upper bounds on the conditional entropy $H(Y|X)$: one coming from the data processing inequality and two derived in a similar way to Fano’s inequality. We defer the proofs to Appendix D for the sake of conciseness, but in broad strokes, our results come from relating the bounds on the error probability provided by these information-theoretic inequalities (typically in the context of error-correcting codes) to the guarantees provided by CP in Theorem 2.1.
@@ 3.1 Data Processing Inequality for Conformal Prediction
We start by using the classical data processing inequality (DPI) in the context of conformal prediction. Specifically, we focus on the DPI for $f$-divergences, which we discuss thoroughly in Appendix C.1. In brief, for a convex function $f$ with $f(1)=0$, the $f$-divergence between two probability measures $P$ and $Q$ is defined as (see [58])
@@
In particular, with $f(x)=x\log x$ we recover the familiar notion of KL-divergence [27]. The DPI for $f$-divergences states that for any two probability measures $P_{X}$ and $Q_{X}$ defined on a space ${\mathcal{X}}$, and any map $W_{Y|X}$, which maps $(P_{X},Q_{X})$ to $(P_{Y},Q_{Y})$, we have
@@
Note that the entropy term $H(Y|X)$ is computed using the measure $P$. We relegate the proof to Appendix D.1. In the bound in (3), the term $Q(Y\in{\mathcal{C}}(X))$ appears inside a log, so an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(X))$ would result in a lower bound and would be biased. We can provide an upper confidence bound on this estimate using the empirical Bernstein inequality [41] and use that instead. Based on the empirical Bernstein inequality, with probability $1-\delta$, we have
@@ 4 Conformal Training
Although split conformal prediction is applicable to any pretrained ML model as a post-processing step, the overall performance of any CP method (commonly measured by its inefficiency) is highly dependent on the underlying model itself. Therefore, previous works have proposed to take CP into account already during model training and directly optimize for low predictive inefficiency [13, 8, 61]. We use the term conformal training to refer to such approaches in general (see Appendix F for an overview of the topic). In particular, we focus on ConfTr [61] since it generalizes and outperforms [8]. In ConfTr [61] each training batch ${\mathcal{B}}$ is split into calibration ${\mathcal{B}}_{cal}$ and test ${\mathcal{B}}_{test}$ halves to simulate the SCP process (see Section 2) for each gradient update of model $f$ and minimize the following size loss
@@
where ${\mathcal{C}}_{f}(x)$ is constructed using the statistics of the nonconformity scores computed on ${\mathcal{B}}_{cal}$. We use the notation ${\mathcal{C}}_{f}(x)$ to emphasize the dependence of the prediction set on model $f$. Still, SCP involves step functions and Stutz et al. [61] introduce two relaxations to recover a differentiable objective: (i) the computation of quantiles is relaxed via differentiable sorting operators [17, 9, 50]; (ii) the thresholding operation in the construction of prediction sets in (2) is replaced by smooth assignments via the logistic sigmoid. The latter relaxation gives “soft” prediction sets $\hat{\mathcal{C}}_{f}(x)$, which contain each of the labels $y\in{\mathcal{Y}}$ with a certain probability. See Algorithm 1 for a depiction of conformal training.
Our upper bounds on $H(Y|X)$, namely DPI, MB Fano and simple Fano, presented in the previous section can be made differentiable in the same way, and thus can also serve as proper loss functions for conformal training. The motivation for doing so is twofold. First, the conditional entropy $H(Y|X)$ captures the underlying uncertainty of the task, or equivalently, the uncertainty under the true labelling distribution $P_{Y|X}$. Thus, by minimizing these upper bounds, we can hope to push the model $f$ closer to the true distribution, which is known to achieve minimal inefficiency [65]. Interestingly, the cross-entropy loss also bounds $H(Y|X)$ from above and thus can be motivated as a conformal training objective from the same angle. In that regard, the DPI bound from Proposition 3.1 is particularly advantageous as it is provably tighter than the cross-entropy (see Appendix D.1).
Second, we can connect the simple Fano bound from Corollary 3.1 to the size loss (7) from [61]. In Appendix F.1, we show that via Jensen’s inequality and $\log(|Y|{-}|C(X)|)\leq\log|Y|$ the bound in (6) can be further upper bounded as
@@
input : A batch of labeled samples ${\mathcal{B}}$, a model $f$, a loss function ${\mathcal{L}}$ that can be any of our upper bounds or other conformal training objectives.
for each training batch ${\mathcal{B}}$ do
       split ${\mathcal{B}}$ into ${\mathcal{B}}_{cal}$ and ${\mathcal{B}}_{test}$
       for each $(x_{i},y_{i})\in{\mathcal{B}}_{cal}$ do
             compute score $s_{i}=s_{f}(x_{i},y_{i})$
      sort scores (in a differentiable manner) obtaining $s_{(1)}<s_{(2)}<\ldots<s_{(|{\mathcal{B}}_{cal}|)}$
       set $\hat{q}=s_{(\lceil(|{\mathcal{B}}_{cal}|+1)(1-\alpha)\rceil)}$
        // get $1-\alpha$ quantile estimate using ${\mathcal{B}}_{cal}$
       for each $(x_{j},y_{j})\in{\mathcal{B}}_{test}$ do
             for each $y\in{\mathcal{Y}}$ do
                   $\hat{\mathcal{C}}_{f}(x_{j},y)\leftarrow\sigma(\hat{q}-s_{f}(x_{j},y))$
                    // Construct soft prediction set
       // The DPI and model-based Fano bounds also require class probabilities under $Q$, which in this case are given by the model $f(x_{j})$
       compute loss according to ${\mathcal{L}}$ on ${\mathcal{B}}_{test}$ using $y_{j},\hat{\mathcal{C}}_{f}(x_{j})$ and if needed $f(x_{j})$
       update $f$ via gradient descent to minimize loss
Algorithm 1 Conformal training algorithm.
@@ 5 Side Information
where $Q_{Z|X,Y}$ is an auxiliary model that predicts the side information given the input and target variables. Such a model could be learned separately from the main model $Q_{Y|X}$ given access to a dataset ${\mathcal{D}}_{side}=\{(x_{i},y_{i},z_{i})\}$. Therefore, we can now calibrate with CP by taking into account the side information and then, at test time, given access to the input and side information, we can use the appropriate probabilities $Q_{Y|X,Z}$ to construct the prediction sets. Intuitively, the prediction sets from such a procedure should be smaller than the prediction sets obtained from using $Q_{Y|X}$ directly. It should be noted that, in the case of side information not being available, we can marginalize $Q_{Y,Z|X}$ over $Z$, which, by construction, falls back to the original model $Q_{Y|X}$. If the availability pattern of $Z$ is consistent between the calibration and test sets, the conformal prediction guarantees still hold by defining the model as
@@
With this addition to the split conformal prediction tool set, if new (side) information is made available at test time, we can properly incorporate it into the CP pipeline without having to retrain the main classifier. One needs only train an auxiliary classifier $Q_{Z|X,Y}$, which might be much simpler than $Q_{Y|X}$ (in our experiments, $Q_{Z|X,Y}$ is given by a single linear layer) and require much less data. One notable example of side information arises naturally in the distributed setting, which we discuss next.
@@ 5.1 The Distributed Learning Setting
This gives us an upper bound on the entropy of the entire population that decomposes into a sum of local functions, one for each client, only requiring local information. Thus, we can easily carry out conformal training for $Q_{Y|X}$ by minimizing this upper bound in the federated setting with, e.g., federated averaging [42]. At test time, we can take the device ID $z$ as side information. To this end, we can either train a model $Q_{Z|X,Y}$ in parallel and use (9) with the global model $Q_{Y|X}$ at test time to get $Q_{Y|X,Z}$, or we can obtain $Q_{Y|X,Z}$ by fine-tuning the global model $Q_{Y|X}$ with local data.
@@ 6 Related Work
Conformal prediction, a powerful framework for uncertainty quantification developed by Vovk and collaborators [64, 60], has recently witnessed a wide adoption in many fields, e.g., healthcare [48, 3, 39, 38] and finance [67, 6]. The marriage of conformal prediction and machine learning has been especially fruitful. Since the seminal work by Vovk et al. [64], many extensions and applications have been proposed, covering topics such as survival analysis [10], treatment effect evaluation [32], classification [21, 4] and regression [54] settings, risk control [5, 7], and covariate shift [62].
To our knowledge, our work represents the first attempt to bridge conformal prediction and information theory. Among other things, this allows us to build on the conformal training ideas of Bellotti [8] and Stutz et al. [61], deriving principled learning objectives that generalize their approaches, dispense with some of their hyperparameters and result in more efficient prediction sets. Further, we empirically show that our conformal training objectives provide a strong enough learning signal to train complex architectures from scratch, with strong results on ResNet-34 and ResNet-50 [23] fitted on CIFAR10 and CIFAR100, respectively. In contrast, the previous state-of-the-art approach, ConfTr, struggles in those settings (see experiments in Section 7) and required pretrained models for consistent results [61]. Further, our information-theoretic interpretation of CP provides a new simple and effective mechanism to leverage side information in split conformal prediction. We are unaware of any other approaches to treat side information within the conformal prediction framework in the literature.
On the information theory side, the notion of $f$-divergence and related inequalities have appeared in many different works. The use of $f$-divergence goes back to works of Ali and Silvey, Csiszár, and Morimoto in the 60s, as in, for instance, [2, 15, 44]. A key $f$-divergence inequality is the data processing inequality—see [58, 57] for an extensive survey. It provides a unified way of obtaining many classical and new results, including Fano’s inequality [19]. The tightness of the data processing inequalities is discussed in terms of Bregman’s divergence in [12, 34] and in terms of $\chi^{2}$-divergence in [57]. List decoding, which is closely connected to CP, was introduced in the context of communication design by Elias [18] and Wozencraft [68]. A generalization of Fano’s inequality to list decoding was given in [16] in the context of multi-user information theory, see also the general Fano inequality for list decoding presented in [53]. Variable-size list decoding was discussed in [57] using ideas first introduced in [52] and [35]. A selection of relevant inequalities for list decoding can be found in [57].